OpenStudy (anonymous):

help me please 9/16 divided by 7/10

OpenStudy (mikeymaximum):

I got 45/56

OpenStudy (anonymous):

thank you but how did you get that I'm confused lol

OpenStudy (mikeymaximum):

kk 1 sec

OpenStudy (anonymous):

@MikeyMaximum Please do not give direct answers. They are against OpenStudy's Code of Conduct. If you keep giving them, you might get suspended. Please refrain from giving direct answers in the future. Thanks. :) If you do not know the rules, then you can read them here: http://openstudy.com/code-of-conduct

OpenStudy (igreen):

^^^^

OpenStudy (igreen):

When we divide fractions, we flip the 2nd fraction and multiply. \(\dfrac{9}{16} \times \dfrac{10}{7} \rightarrow \dfrac{9 \times 10}{16 \times 7}\) Multiply 9 * 10 and 16 * 7 @NadiaElhaj_1

OpenStudy (mikeymaximum):

I SAID "kk 1 sec" WHEN SHE ASKED FOR AN EXPLANATION... -_- R u serious @iGreen

OpenStudy (igreen):

@MikeyMaximum We have to GUIDE the asker to the answer, we need her PARTICIPATION, otherwise it is a direct answer.

OpenStudy (igreen):

You can't do all the work and give the answer..otherwise the asker can just take the answer and doesn't learn how to do it..

OpenStudy (anonymous):

thank you @iGreen and @MikeyMaximum thank you so much :)